OpenStudy (shes.radical):

Given the following functions f(x) and g(x), solve (f ⋅ g)(3) and select the correct answer below: f(x) = 4x2 + 12 g(x) = x − 1 (a) 47 (b) -47 (c) 96 (d) -96

OpenStudy (shes.radical):

\[f(x)=4x^2+12\] \[g(x)=x-1\]

undeadknight26 (undeadknight26):

Sorry m8 I hate these.

OpenStudy (shes.radical):

@Vocaloid

OpenStudy (shes.radical):

@Mehek14

OpenStudy (shes.radical):

@just_one_last_goodbye

OpenStudy (just_one_last_goodbye):

I dont know either ask @nincompoop or @Squirrels

OpenStudy (squirrels):

add me on myspace, snapchat, kik, friendster, instagram, twitter, facebook, and tumblr @Yung_Squirrels

OpenStudy (shes.radical):

if u answer the question -.-

OpenStudy (squirrels):

UGHH ok fine lol one second

OpenStudy (shes.radical):

@twistnflip

OpenStudy (anonymous):

(4x^2 + 12)(x-1) FOIL can you do that?

OpenStudy (squirrels):

thanks homie @twistnflip i needed to go anyway. its been gucci. peace.

OpenStudy (shes.radical):

i got 312..

OpenStudy (anonymous):

Well that is obviously wrong, so... Please do what I said you should get a quadratic once simplified

OpenStudy (shes.radical):

i did do what you said..and somehow i got 312.. 4(3)^2=144+12=156*2=312..

OpenStudy (anonymous):

Actually just plug 3 in here: (4(3)^2 + 12)((3)-1) (4 * 9 + 12) (2) (48)(2) = 96

OpenStudy (shes.radical):

...wow. thanks so much xD

OpenStudy (dinamix):

its 96 i find it also